Z varies directly as the square of x and inversely as the square root of y. Find the percentage change in z if x increases by 20% and y decreases by 19%.
Answer by Boreal(15235)

z=kx^2*y^(-1/2)
z new=k(1.2x)^2*(0.81y)^(-1/2)
The ratio has the k cancel, and it is 1.44/(1/0.9), for 1.2^2=1.44 and 1/sqrt(0.81)=1/0.9.
That is 1.44*0.9=1.296
29.6% increase.